Calculate National income by income method and expenditure method:
Items Rs in Crores
Subsidies 5
Private Final consumption expenditure 100
NFIA -10
Indirect taxes 25
Rent 5
Government Final consumption expenditure 20
Net domestic fixe capital formation 30
Operating surplus 20
Wages 50
Net exports -5
Addition to stock -5
Social security contribution by employers 10
Mixed income 40
(
1
Expert's answer
2020-06-02T08:40:40-0400

Expenditure method:

National Income (NI) = C+I+G+(X-M)

National Income (NI) = Private Final consumption expenditure 100 + Government Final consumption expenditure 20 + (Net domestic fixed capital formation 30 + Addition to stock -5) + Net exports -5

National Income (NI) = 100 + 20 + 30-5-5

National Income (NI) = 140


Income method

National Income (NI) = Rent + Wages + Operating surplus + Social security contribution by employers + Mixed income + (Indirect taxes - Subsidies)

National Income (NI) = Rent 5 + Wages 50 + Operating surplus 20 + Social security contribution by employers 10 + Mixed income 40 + (Indirect taxes 25 - Subsidies 5)

National Income (NI) = 5 + 50 + 20 + 10 + 40+25-5

National Income (NI) = 145